OpenStudy (anonymous):

Help with adding rational expressions. How do I add 5+ sqrt(x) + (1/1+x)

OpenStudy (anonymous):

\[5+ \sqrt(x) + (1 \div (1+x))\]

Nnesha (nnesha):

\[\huge\rm 5+\sqrt{x} +\frac{1}{1+x}\] find the common denominator

Nnesha (nnesha):

5+sqrt{x} is same as 5+sqrt{x} over one \[\huge\rm \frac{5+\sqrt{x}}{1} +\frac{1}{1+x}\] so what is the common denominator

OpenStudy (anonymous):

1+x?

Nnesha (nnesha):

yes right so when we find common denominator we should multiply the numerator of 1st fraction by the denominator of 2nd fraction and multiply the numerator of 2nd fraction by the denominator of 1st fraction (just like cross multiply\[\huge\rm \frac{ (1+x)(5+\sqrt{x} )+1(1)}{ 1+x }\] )